OpenStudy (anonymous):

Two automobiles start at the same time and place, and travel in opposite directions. The first car averages 32 km an hour and the second, 36 km an hour. In how many hours will the cars be 306 km. apart.

OpenStudy (anonymous):

Please do help me..

OpenStudy (hoblos):

32x + 36x = 306 68x=306 x= 4.5 h

OpenStudy (anonymous):

32x + 36x = 306

OpenStudy (anonymous):

thank you so much..

OpenStudy (hoblos):

my pleasure XD

OpenStudy (anonymous):

4 1/2 hrs = 4 hrs 30 mins.